# A geometric perspective on Algebraic Quantum Field Theory

### v1

@@ Abstract
In this paper we give a streamlined overview of some of the recent constructions provided with K.-H. Neeb, G. Ólafsson and collaborators for a new geometric approach to Algebraic Quantum Field Theory (AQFT). Motivations, fundamental concepts and some of the relevant results about the abstract structure of these models are here presented.
@@ 1 Introduction
These notes aim to provide a structured guideline, building upon recent results developed in collaboration with K.-H. Neeb and G. Ólafsson, to explore a geometric perspective on Algebraic Quantum Field Theory. In this presentation, we will focus on some selected results from the literature of recent years, illustrating them through specific cases. We hope this approach will assist readers in navigating the main findings and the research directions, while maintaining a balanced view of both the geometric and operator-algebraic perspectives on the related topics.
@@
We start our discussion with some motivations. Algebraic Quantum Field Theory (AQFT) is an axiomatic framework for Quantum Field Theory where a model is specified by a map associating to any open region of the spacetime its von Neumann algebra of local observables, acting on a fixed complex Hilbert space $\mathcal{H}$ (the state space), satisfying fundamental quantum and relativistic assumptions as isotony, locality, Poincaré covariance, positivity of the energy and cyclicity of the vacuum vector for local algebras [Ha96]. In this picture a rich interplay between the geometry and the algebraic parts of the theory appears.
@@
The algebraic structure of the AQFT models, through Tomita–Takesaki theory, contains a large part of the information on the geometry of the model. The Bisognano-Wichmann property and the PCT theorem describe a geometric action of the Tomita modular group and the Tomita modular conjugation of local algebras of some specific regions with respect to the vacuum state. They have been verified for a large number of models (see e.g. [BW75, BGL93, Mu01, Mo18, DM20]) and this ensures a very strong relation between the geometry and the algebraic structure of the models that enriches in many directions in the Algebraic approach to Quantum Field Theory. As one of the examples of this interplay we point out that recently operator algebraic methods have proven to be highly effective in studying energy conditions of AQFT models. Here the modular Hamiltonian, namely the logarithm of the Tomita modular operator associated with the local algebra of a specific wedge region with respect to a specific state, plays a central role, see for instance [MTW22, Lo20, CLRR22, CLR20, Ara76, LX18, LM23, LM24, Wi18]. Indeed, by the Bisognano–Wichmann Theorem it has a geometrical meaning and a rigorous analysis from both the algebraic and the geometric perspectives is possible.
@@
Consider an AQFT on Minkowski spacetime. A Rindler wedge and its Poincaré transforms (wedge regions) are fundamental localization regions for von Neumann algebras or particle states and are in 1-1 correspondence with a one–parameter group of boost symmetries (properly parametrized) that fix them as a subset of Minkowski spacetime. The algebraic canonical construction of the free field provided by Brunetti–Guido–Longo (BGL) builds on the wedge-boost correspondence, the Bisognano–Wichmann (BW) property and the PCT Theorem (or the geometric action of the modular conjugation, here called modular reflection), cf. [BGL02]. This approach is actually very general and this is already visible in the BGL paper. Indeed the authors applied this construction to many of the free models of physical interests that have a 1-1 correspondence between specific regions of spacetime and one–parameter subgroups of the symmetry group: In free (possibly conformal) theories on Minkowksi spacetime, on de Sitter space and on the chiral circle.
What we learn is that we can associate bijectively wedge regions of the spacetime to one–parameter subgroups of the Lie group, hence to generators in the Lie algebra. This is the starting point of our analysis. The core of this analysis relies on the understanding of a deep connection between the geometry of real subspaces (the language for localized particle states), given by the Tomita modular operator and modular conjugation, and the geometry of specific elements in the Lie algebra of a Lie group $G$, called Euler elements. We call an element $x$ of the finite dimensional real Lie algebra ${\mathfrak{g}}$ an Euler element if $\mathop{{\rm ad}}\nolimits x$ is diagonalizable with ${\rm Spec}(\mathop{{\rm ad}}\nolimits x)\subseteq\{-1,0,1\}$, hence $\mathop{{\rm ad}}\nolimits x$ defines a $3$-grading of ${\mathfrak{g}}$:
${\mathfrak{g}}={\mathfrak{g}}_{1}(x)\oplus{\mathfrak{g}}_{0}(x)\oplus{\mathfrak{g}}_{-1}(x),\quad\mbox{ where }\quad{\mathfrak{g}}_{\nu}(x)=\ker(\mathop{{\rm ad}}\nolimits x-\nu\mathop{{\rm id}}\nolimits_{\mathfrak{g}})$
(see [BN04, Ka00] for more details on Euler elements and three graded Lie algebras). The combination of the language of standard subspaces and Euler elements is useful for several reasons. In particular it provides a general structure that matches the geometry of the models with Tomita modular theory and it captures the fundamental localization properties of particle states and von Neumann algebras of observables. This allows to define an abstract framework that properly generalizes the set of AQFT models.
@@
We start this discussion by recalling some fundamental models in Algebraic Quantum Field Theory and introducing our geometric perspective. Once the correspondence between wedges and boosts is clear, we refer it to the notion of abstract wedges, their properties and a classification of simple real Lie algebras supporting such abstract wedges. Then we will see how models in AQFT can be generalized when abstract wedges are taken into account. This first part mainly refers to [MN21]. The fundamental models we are discussing and the related axiomatic framework sits on a causal homogeneous space [FNÓ23, MNO23a, MNO23b, NÓØ21, NÓ21, NÓ22]. Then we pass to discuss a recent construction of non-modular covariant net of standard subspaces, cf. [MN22, Sect. 3.1]. In Sect. 3 we introduce the Euler Element Theorem that ensures that it is possible to deduce our Lie theoretical description of wedges from the Bisognano–Wichmann property and a regularity property. The latter property is strongly related with a localization property, see Sects. 3.2 and 3.3 and [MN24, Sect. 4]. Lastly we show that the von Neumann algebras appearing in this picture are of type III1 (the Connes classification), see Sect 3.4 and [MN24, Sect. 5]. Many of these results are formulated for general von Neumann algebras and/or standard subspaces in some specific relative position. We explain how they apply to nets of von Neumann algebras (resp. standard subspaces). This approach provides feedback for representation theory (as new constructions of Lie group representations as in [FNÓ23, NÓ21]) and for the algebraic approach to Quantum Field Theory without restrictions to second quantization models. This paper will mainly deal with abstract wedges, a more detailed overview on the properties of wedge regions in causal symmetric space is postponed.
@@
Acknowledgment: We acknowledge Gerardo Morsella for stimulating discussions and Karl-Hermann Neeb for useful comments. We thank the MIUR Excellence Department Project 2023-2027 MatMod@TOV, CUP E83C23000330006, awarded to the Department of Mathematics of University of Rome Tor Vergata. We also thanks INdAM-GNAMPA.
@@
  • Unitary or (anti-)unitary representations on the complex Hilbert space $\mathcal{H}$ are denoted as pairs $(U,\mathcal{H})$. For a unitary representation $(U,\mathcal{H})$ of $G$ we write: $\partial U(x)=\frac{d}{dt}\big{|}_{t=0}U(\exp tx)$ for the infinitesimal generator of the unitary one–parameter group $(U(\exp tx))_{t\in{\mathbb{R}}}$ in the sense of Stone’s Theorem.
  • If $G$ is a group acting on a set $M$ and $W\subseteq M$ a subset, then the stabilizer subgroup of $W$ in $G$ is denoted $G_{W}:=\{g\in G\colon g.W=W\}$, and $S_{W}:=\{g\in G\colon g.W\subseteq W\}$.
@@ 2 Preliminaries
2.1 Wedge regions in spactime
@@
$\mathcal{O}^{\prime}$ represents the set of points that correspond to events that do not have causal dependence form events in the region $\mathcal{O}$. The connected component of the identity of the homogenous symmetry group of $M$ is the proper orthochronous Lorentz group $\mathcal{L}_{+}^{\uparrow}=\mathop{{\rm SO}}\nolimits(1,d)_{e}$. We denote by $\mathcal{L}_{+}$ the proper Lorentz group, namely the group generated by $\mathcal{L}_{+}^{\uparrow}$ and the space and time reflection
@@
acting on the first two coordinates, hence $\mathcal{L}_{+}=\mathcal{L}_{+}^{\uparrow}\cup\mathcal{L}_{+}^{\downarrow}$ where $\mathcal{L}_{+}^{\downarrow}=j_{1}\mathcal{L}_{+}^{\uparrow}$. The associated inhomogeneous symmetry groups, will be denoted by $\mathcal{P}_{+}^{\uparrow}={\mathbb{R}}^{1,d}\rtimes\mathcal{L}_{+}^{\uparrow}$ and $\mathcal{P}_{+}={\mathbb{R}}^{1,d}\rtimes\mathcal{L}_{+}$, the proper orthochronous Poincaré group and the proper Poincaré group, respectively, and $\mathcal{P}_{+}=\mathcal{P}_{+}^{\uparrow}\cup\mathcal{P}_{+}^{\downarrow}=\mathcal{P}_{+}^{\uparrow}\cup j_{1}\mathcal{P}_{+}^{\uparrow}$.
@@
In this framework there are some distinguished regions (connected open subsets) of $M$ that have a deep connection with the symmetry group. A wedge region $W\subset{\mathbb{R}}^{1,d}$ is a Poincaré transform of the standard right wedge $W_{R}$:
$W=gW_{R},\quad g\in\mathcal{P}_{+}^{\uparrow},\quad\text{where}\quad W_{R}=\left\{x\in{\mathbb{R}}^{1+d}:|x_{0}|<x_{1}\right\}$
Now, consider the one–parameter group of boosts contained in $\mathcal{P}_{+}^{\uparrow}$:
$\Lambda_{W_{R}}(t)=\left(\begin{array}[]{ccc}\cosh(t)&\sinh(t)&\textbf{0}\\ \sinh(t)&\cosh(t)&\textbf{0}\\ \textbf{0}&\textbf{0}&\textbf{1}\end{array}\right)$ (1)
All the other Poincaré boosts are determined by the adjoint action $\Lambda_{W}=g\Lambda_{W_{R}}g^{-1}$, with $g\in\mathcal{P}_{+}^{\uparrow}$. It is not by coincidence that we label the boost with the same index of a wedge region. Indeed $\Lambda_{W}(t)W=W$ and, once the parametrization is fixed, the 1-1 correspondence
$W\stackrel{{\scriptstyle 1-1}}{{\longleftrightarrow}}\Lambda_{W}.$
Let $h_{W}\in\mathop{\bf L{}}\nolimits(\mathcal{P}_{+}^{\uparrow})$ be the generator of $\Lambda_{W}(t)=\exp(th_{W})$. Then there is a 1-1 correspondence between wedge regions and the Lie algebra elements that are generators of boost transformations
$h_{W}\stackrel{{\scriptstyle 1:1}}{{\longleftrightarrow}}\Lambda_{W}(t)=\exp(th_{W})\stackrel{{\scriptstyle 1:1}}{{\longleftrightarrow}}W.$
@@
Let $W=gW_{R}$ and $W^{\prime}=\{x\in M:(x-y)^{2}<0,y\in W\}^{\circ}$ be the causal complement wedge of a wedge $W$, the transformation $j_{W}:=gj_{W_{R}}g^{-1}\in\mathcal{P}_{+}^{\mathop{\downarrow}}$ maps $W$ onto $W^{\prime}$. Clearly $j_{W}=j_{W^{\prime}}$ and it is important to remark that $j_{W}=\exp(i\pi h_{W})$. We have learnt that wedge regions can be determined by Lie algebra elements and that there is a natural transformation sending $W$ to $W^{\prime}$ that happens to be the analytic extension of the one–parameter group $\Lambda_{W}$. We further remark that $\Lambda_{W^{\prime}}(t)=\Lambda_{W}(-t)$, hence $h_{W^{\prime}}=-h_{W}$. When conformal symmetries are considered, namely the symmetry group is the group of local diffeomorphisms (defined out of meager sets) which preserve the metric tensor up to a positive function, then all the conformal transforms of a wedge in Minkowski space, such as wedges, double cones, future cones and past cones, give rise to conformal wedges and the previous identification with Lie generators continues to hold, see [BGL02, BGL93]. Here, a future and a past cone are intended to be respectively regions of the form $(C+a)^{\circ}$ and $(-C+b)^{\circ}$ for some $a,b\in M$ and a double cone is a region of the form $(C+a)^{\circ}\cap(-C+b)^{\circ}$ where $b-a\in C^{\circ}$.
@@
We will see now that these examples of this interesting correspondence between wedges and one–parameter groups of boost generators are special cases of a more general picture.
2.2 Euler elements

### v2

@@ Abstract
In this paper we give a streamlined overview of some of the recent constructions provided with K.-H. Neeb, G. Ólafsson and collaborators for a new geometric approach to Algebraic Quantum Field Theory. Motivations, fundamental concepts and some of the relevant results about the abstract structure of these models are here presented.
@@ 1 Introduction
These notes aim to provide a structured guideline, building upon recent results developed in collaboration with K.-H. Neeb and G. Ólafsson, to explore a geometric perspective on Algebraic Quantum Field Theory. In this presentation, we will focus on some selected outcomes from the literature of recent years, illustrating them through specific cases. We hope this approach will assist readers in navigating the main findings and the research directions, while maintaining a balanced view of both the geometric and operator-algebraic perspectives on the related topics.
@@
We start our discussion with some motivations. Algebraic Quantum Field Theory (AQFT) is an axiomatic framework for Quantum Field Theory where a model is specified in the sense of the Haag-Kastler net by a map associating to any open region $\mathcal{O}$ of the spacetime $M$ its von Neumann algebra $\mathcal{A}(\mathcal{O})$ representing the set of local observables, acting on a fixed complex Hilbert space $\mathcal{H}$ (the state space), satisfying fundamental quantum and relativistic assumptions as:
  • Isotony: $\mathcal{O}_{1}\subseteq\mathcal{O}_{2}$ implies $\mathcal{A}(\mathcal{O}_{1})\subseteq\mathcal{A}(\mathcal{O}_{2})$ ;
  • Locality: $\mathcal{O}_{1}\subseteq\mathcal{O}_{2}^{\prime}$ implies $\mathcal{A}(\mathcal{O}_{1})\subseteq\mathcal{A}(\mathcal{O}_{2})^{\prime}$, where $\mathcal{O}^{\prime}$ is the “causal complement” of $\mathcal{O}$, i.e., the maximal open subset that cannot be connected to $\mathcal{O}$ by causal curves;
  • Covariance: There is a Lie group $G$ acting on $M$ and a unitary representation $U\colon G\to\mathop{\rm U{}}\nolimits(\mathcal{H})$ such that $U_{g}\mathcal{A}(\mathcal{O})U_{g}^{-1}=\mathcal{A}(g\mathcal{O})$ for $g\in G$;
  • Uniqueness and Invariance of vacuum: there exists a unique (up to a phase) unit vector $\Omega\in\mathcal{H}$ such that $U(g)\Omega=\Omega$ for every $g\in G$. $\Omega$ is called the vacuum state;
  • Reeh–Schlieder property: $\Omega\in\mathcal{H}$ is cyclic for $\mathcal{A}(\mathcal{O})$ if $\mathcal{O}\not=\emptyset$
(cf. [Ha96]). Note that given an open region $\mathcal{O}_{1}$, if there exists an open region $\mathcal{O}_{2}\subset\mathcal{O}_{1}^{\prime}$ then, by locality and Reeh-Schlieder property, $\Omega$ is also separating for $\mathcal{A}(O_{1})$. In this framework there is a deep interplay between the geometric and the algebraic parts of the theory.
The Tomita-Takesaki theory states that given a von Neumann algebra $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$ with a cyclic and separating vector $\Omega\in\mathcal{H}$, one can define the Tomita operator $S_{\mathcal{A},\Omega}$ as the closure of the anti-linear involution
The polar decomposition $S_{\mathcal{A},\Omega}=J_{\mathcal{A},\Omega}\Delta^{\frac{1}{2}}_{\mathcal{A},\Omega}$ defines an antiunitary involution $J_{\mathcal{A},\Omega}$ called modular involution and a self-adjoint positive operator $\Delta_{\mathcal{A},\Omega}$ called modular operator satisfying
$J_{\mathcal{A},\Omega}\mathcal{A}J_{\mathcal{A},\Omega}=\mathcal{A}^{\prime},\quad\Delta_{\mathcal{A},\Omega}^{it}\mathcal{A}\Delta_{\mathcal{A},\Omega}^{-it}=\mathcal{A},\quad J_{\mathcal{A},\Omega}\Delta_{\mathcal{A},\Omega}J_{\mathcal{A},\Omega}=\Delta_{\mathcal{A},\Omega}^{-1},$
The one–parameter group $\Delta_{\mathcal{A},\Omega}^{it}$ is called the modular group. It is relevant to remark that couples of an antiunitary operator $J$ and a self-adjoint positive operator $\Delta$ on a complex Hilbert space $\mathcal{H}$ satisfying $J\Delta J=\Delta^{-1}$ are in 1-1 correspondence with standard subspaces ${\sf H}\subset\mathcal{H}$, namely real closed subspaces ${\sf H}\subset\mathcal{H}$ such that $\overline{{\sf H}+i{\sf H}}=\mathcal{H}$ and ${\sf H}\cap i{\sf H}=\{0\}$, through the condition
${\sf H}=\ker(1-S)$ (1)
where $S=J\Delta^{\frac{1}{2}}$ (cf. Sect. 2.4.2). Because of this relation with the Tomita operator, the language of standard subspaces has a key role in our analysis.
Now, consider a net of von Neumann algebras $\mathcal{A}$ on the $1+d$ dimensional Minkowksi space ${\mathbb{R}}^{1,d}$. Let $\mathcal{A}(W_{R})$ be von Neumann algebra of observables associated to the Rindler wedge
$W_{R}:=\{x=(x_{0},\ldots,x_{d})\in{\mathbb{R}}^{1,d}:|x_{0}|<x_{1}\},$ (2)
The Bisognano-Wichmann property states that the modular group of $\mathcal{A}(W_{R})$ with respect to the vacuum state $\Omega$ unitarily implements the following one–parameter subgroup of the Poincaré group, known as Lorentz boosts:
$\Lambda_{W_{R}}(t)=\left(\begin{array}[]{ccc}\cosh(t)&\sinh(t)&\textbf{0}\\ \sinh(t)&\cosh(t)&\textbf{0}\\ \textbf{0}&\textbf{0}&\textbf{1}\end{array}\right).$ (3)
Specifically, this implementation is given by
$\Delta_{\mathcal{A}(W_{R}),\Omega}^{it}=U(\Lambda_{W_{R}}(-2\pi t)).$
The modular conjugation also has a geometric action when the modular reflection property is satisfied:
$U(j_{W_{R}})=J_{\mathcal{A}(W),\Omega}$
@@
where $j_{W_{R}}=\mathop{{\rm diag}}\nolimits(-1,-1,\textbf{1}_{d-1})$ reflects $j_{W_{R}}W_{R}=W_{R}^{\prime}$. It is evident that the Bisognano-Wichmann property together with the modular reflection property (see also the PCT Theorem, Section 2.4.1) determine a geometric action of the modular group and the modular conjugation of the local algebras associated with certain specific regions. They have been verified for a large number of models (see e.g. [BW75, BGL93, Mu01, Mo18, DM20]), ensuring a very strong relation between the geometry and the algebraic structures, which enriches the algebraic approach to Quantum Field Theory in many directions. For instance, in the recent years, operator algebraic methods have proven to be highly effective in studying energy conditions of AQFT models. Here the modular Hamiltonian, namely the logarithm of the modular operator, associated with the local algebra of wedge regions plays a central role, see for instance [MTW22, Lo20, CLRR22, CLR20, Ara76, LX18, LM23, LM24, Wi18]. By the Bisognano–Wichmann property it has a geometrical meaning and a rigorous analysis from both the algebraic and the geometric perspectives is possible.
@@
One can establish a 1-1 correspondence between the Poincaré transforms of the Rindler wedge called wedge regions $W\subset{\mathbb{R}}^{1,d}$ and one–parameter group of boosts $\Lambda_{W}$ obtained by Poincaré adjoint action on $\Lambda_{W_{R}}$. The algebraic canonical construction of the fundamental model called the free field provided by Brunetti–Guido–Longo (BGL) builds on this wedge-boost correspondence, the Bisognano–Wichmann (BW) property and the modular reflection property, cf. [BGL02]. Here, the standard subspace techniques play a prominent role. Given a (anti-)unitary representation $U$ of the Poincaré group (a particle in QFT) the localization properties of particle states (vectors of the Hilbert space $\mathcal{H}$ supporting $U$) on wedge regions can be characterized using nets of standard subspaces: There is a map associating to each wedge region $W$ a standard subspace ${\sf H}(W)\subset\mathcal{H}$ of vector states in a canonical way by assuming the Bisognano-Wichmann property, modular reflection property, and taking into account the correspondence between the Tomita operators and standard subspaces given by (1). The free net of von Neumann algebras is then constructed by a canonical procedure called second quantization. This approach is actually very general and this is already visible in the BGL paper. Indeed, the authors apply this construction to many of the free models of physical interest that have an analogue 1-1 correspondence between specific regions of spacetime and one–parameter subgroups of the symmetry group: In free (possibly conformal) theories on Minkowksi spacetime, on de Sitter space and on the chiral circle (see also [GLW98, Lo08]).
In this work we present a geometric perspective on these examples that strictly generalizes the framework given by the AQFT. The paper starts giving more details on the wedge-boost generator correspondence in some of the AQFT spacetime (Sect. 2.1). We will explore how this picture can be generalized using a geometric approach, following [MN21]. The Lie algebra elements generating the one-parameter groups of boosts that identify wedge regions are Euler elements: Let $x$ be an element of ${\mathfrak{g}}$ a finite dimensional real Lie algebra, $x$ is an Euler element if $\mathop{{\rm ad}}\nolimits x$ is diagonalizable with ${\rm Spec}(\mathop{{\rm ad}}\nolimits x)\subseteq\{-1,0,1\}$ (see [BN04, Ka00] for more details on Euler elements and three graded Lie algebras). In Section 2.2 we will see how this definition leads to a notion of abstract wedges, called Euler wedges, for a $\mathbb{Z}_{2}$-graded Lie group whose Lie algebra contains Euler elements without necessity to refer to a specific manifold. Simple real Lie algebras that support Euler elements, can be classified [MN21, Thm. 3.10]. One can study causal homogenous spaces to recover a manifold playing the role of the spacetime. In Section 2.3 we present a suitable definition of wedge regions on causal homogeneous spaces and as an example we will see how the de Sitter spacetime can be recovered in this general picture (cf. [NÓ23, NÓ22, MNO23a, MNO23b]).
@@
This allows us to exhibit a very general setting for nets of standard subspaces and von Neumann algebras having as an index set the Euler wedges. We will see in Section 2.4 that this construction is possible for a large family of Lie groups and provides several new models. An axiomatic framework is also provided for nets of standard subspaces on homogeneous spaces see [FNÓ23, MNO23a, MNO23b, NÓØ21, NÓ21, NÓ22]. In Section 3.1 we discuss a recent construction of non-modular covariant net of standard subspaces, namely nets of standard subspaces with no geometric action of the wedge subspace modular groups as in [MN22, Sect. 3.1]. In Section 3.2 we introduce the Euler Element Theorem that ensures that it is possible to deduce the keyrole of Euler elements and Euler wedges from the Bisognano–Wichmann property and a regularity property. The latter property is strongly related with a localization property, as reported in sections 3.2 and 3.3 and in [MN24, Sect. 4]. Lastly we show that the von Neumann algebras appearing in this picture are of type III1 for the Connes classification, see Section 3.4 and [MN24, Sect. 5]. Many of these results are formulated for general von Neumann algebras and/or standard subspaces in some specific relative position. We explain how they apply to nets of standard subspaces and of von Neumann algebras. This approach provides feedback for representation theory (see e.g. constructions of nets of standard subspaces on homogeneous spaces in [FNÓ23, NÓ21]) and for the algebraic approach to Quantum Field Theory without restrictions to second quantization models. This paper will mainly deal with abstract wedges, a more detailed overview on the properties of wedge regions in causal symmetric space is postponed.
@@
Acknowledgments: V.M. thanks Gerardo Morsella for stimulating discussions and Karl-Hermann Neeb for useful comments. V.M. acknowledges the support of the University of Rome Tor Vergata funding OANGQS CUP E83C25000580005, the MUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C23000330006 and INdAM-GNAMPA.
@@
  • $\mathcal{C}_{0}^{\infty}(M)$ denotes the set of compactly supported functions in $M$.
@@
  • Unitary or (anti-)unitary representations on the complex Hilbert space $\mathcal{H}$ are denoted as pairs $(U,\mathcal{H})$. For a unitary representation $(U,\mathcal{H})$ of a Lie group $G$ we write: $\partial U(x)=\frac{d}{dt}\big{|}_{t=0}U(\exp tx)$ for the infinitesimal generator of the unitary one–parameter group $(U(\exp tx))_{t\in{\mathbb{R}}}$ in the sense of Stone’s Theorem. In this setting $C_{U}:=\{x\in{\mathfrak{g}}\colon-i\partial U(x)\geq 0\}$.
@@
  • ${\tt d}U\colon{\mathfrak{g}}\to\mathop{{\rm End}}\nolimits(\mathcal{H}^{\infty})$ for the representation of the Lie algebra ${\mathfrak{g}}$ on the space $\mathcal{H}^{\infty}$ of smooth vectors. Then $\partial U(x)=\overline{{\tt d}U(x)}$ (operator closure) for $x\in{\mathfrak{g}}$.
@@ 2 Preliminaries
In this section we introduce the notations, definitions, theorems and the general techniques that will be employed throughout the paper. We will start by examining the geometry of fundamental examples in Section 2.1. From them it will emerge the key role of Euler elements and the structure of Euler wedges as explained in Section 2.2. Wedge domains in causal homogeneous spaces are briefly recalled in Section 2.3. In Sect. 2.4 we present the fundamental free field construction (Sect. 2.4.1) and we introduce the standard subspace formalism (Sect. 2.4.2) to focus on the axiomatic framework for nets of standard subspaces on abstract wedges and on the generalized BGL-construction provided in [MN21] (Sect. 2.4.3).
2.1 Wedge regions in some spacetimes
@@
$\mathcal{O}^{\prime}$ represents the set of points that are causally independent of those in $\mathcal{O}$. A causally complete region satisfies the property $\mathcal{O}=\mathcal{O}^{\prime\prime}$. The connected component of the identity of the homogenous symmetry group of $M$ is the proper orthochronous Lorentz group $\mathcal{L}_{+}^{\uparrow}=\mathop{{\rm SO}}\nolimits(1,d)_{e}$. We denote by $\mathcal{L}_{+}$ the proper Lorentz group, namely the group generated by $\mathcal{L}_{+}^{\uparrow}$ and the space and time reflection
@@
acting on the first two coordinates, hence $\mathcal{L}_{+}=\mathcal{L}_{+}^{\uparrow}\cup\mathcal{L}_{+}^{\downarrow}$ where $\mathcal{L}_{+}^{\downarrow}=j_{1}\mathcal{L}_{+}^{\uparrow}$. The associated inhomogeneous symmetry groups denoted by $\mathcal{P}_{+}^{\uparrow}={\mathbb{R}}^{1,d}\rtimes\mathcal{L}_{+}^{\uparrow}$ and $\mathcal{P}_{+}={\mathbb{R}}^{1,d}\rtimes\mathcal{L}_{+}$ are respectively the proper orthochronous Poincaré group and the proper Poincaré group. In particular $\mathcal{P}_{+}=\mathcal{P}_{+}^{\uparrow}\cup\mathcal{P}_{+}^{\downarrow}=\mathcal{P}_{+}^{\uparrow}\cup j_{1}\mathcal{P}_{+}^{\uparrow}$.
@@
In this framework there are some distinguished regions (connected open subsets) of $M$ that have a deep connection with the symmetry group. A wedge region
$W=gW_{R}\subset{\mathbb{R}}^{1,d}\quad g\in\mathcal{P}_{+}^{\uparrow}$
is a Poincaré transform of the Rindler wedge $W_{R}$ (here also called the right wedge) introduced in (2).
Now, consider the one–parameter group of boosts $\Lambda_{W_{R}}(t)$ contained in $\mathcal{P}_{+}^{\uparrow}$, as in (3). All the other one-parameter group of boosts are determined by the Poincaré adjoint action $\Lambda_{W}=g\Lambda_{W_{R}}g^{-1}$, with $g\in\mathcal{P}_{+}^{\uparrow}$. There is a 1-1 correspondence between wedge regions and one-parameter group of boosts and their generators as
$h_{W}\stackrel{{\scriptstyle 1:1}}{{\longleftrightarrow}}\Lambda_{W}=g\Lambda_{W_{R}}g^{-1}\stackrel{{\scriptstyle 1:1}}{{\longleftrightarrow}}W=gW,\qquad g\in\mathcal{P}_{+}^{\uparrow},$
where $h_{W}\in\mathop{\bf L{}}\nolimits(\mathcal{P}_{+}^{\uparrow})$ generates $\Lambda_{W}(t)=\exp(th_{W})$.
@@
Let $W=gW_{R}$ and $W^{\prime}=\{x\in M:(x-y)^{2}<0,y\in W\}^{\circ}$, the transformation $j_{W}:=gj_{W_{R}}g^{-1}\in\mathcal{P}_{+}^{\mathop{\downarrow}}$ maps $W$ onto $W^{\prime}$. Clearly $j_{W}=j_{W^{\prime}}$ and it is important to remark that $j_{W}=\exp(i\pi h_{W})$. We have observed that the wedge regions can be characterized by Lie algebra elements and that there exists a natural transformation mapping $W$ to $W^{\prime}$, which coincides with the analytic extension of the one-parameter group $\Lambda_{W}(t)$. We further remark that $\Lambda_{W^{\prime}}(t)=\Lambda_{W}(-t)$, hence $h_{W^{\prime}}=-h_{W}$. When conformal symmetries are considered - specifically, when the symmetry group consists of local diffeomorphisms (defined outside meager sets) that preserve the metric tensor up to a positive function - all conformal transformations of a wedge in Minkowski space, such as wedges, double cones, future cones, and past cones, are referred to as conformal wedge, see [BGL02, BGL93]. Here, a future and a past cone are intended to be respectively regions of the form $(C+a)^{\circ}$ and $(-C+b)^{\circ}$ for some $a,b\in M$ and a double cone is a region of the form $(C+a)^{\circ}\cap(-C+b)^{\circ}$ where $b-a\in C^{\circ}$. The identification of conformal wedges with conformal Lie generators remains valid.
@@
We will now observe that these examples of correspondences between wedges and one-parameter groups of boost generators are special cases of a more general framework.
2.2 Euler elements and abstract wedges